OpenStudy (anonymous):

Find the midpoint of the line segment joining (3, 5) and (-1, 6). help please

ganeshie8 (ganeshie8):

use the midpoint formula

OpenStudy (anonymous):

mid pint of x1,y1 and x2,y2 is X=(x1+x2)/2 and Y=(y1+y2)/2

OpenStudy (anonymous):

@muppz can u do it nw

OpenStudy (anonymous):

so its

OpenStudy (anonymous):

\[-\frac{ -1+3 }{ 2 }\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[\frac{ 6+5 }{ 2}\]

OpenStudy (anonymous):

simply use the midpoint formula (x1+x2)/2,(y1+y2)/2 oin this case it is (3-1)/2,(5+6)/2 =1,11/2

OpenStudy (anonymous):

so 1,5.5

OpenStudy (anonymous):

is that my anaser?

ganeshie8 (ganeshie8):

(1, 5.5) = \(\checkmark\)

ganeshie8 (ganeshie8):

good work ;)